# Adaptive Step Duration for Precise Foot Placement: Achieving Robust Bipedal Locomotion on Terrains with Restricted Footholds

### v2

@@ III Swing Foot Placement Planning via Discrete Model Predictive Control
The main contribution of this work is to design a foot placement planner on the stepping stones with high accuracy and resistance to external perturbations. The stepping stones can be modeled as a sequence of bounds on each step position, which leads to significant constraints on the dynamic walking. Therefore, multiple future steps must be considered to find a stable walking gait by exploiting the viable evolution of the initial DCM with variable step duration.
To simplify the denotation, we define $\mathbf{z}^{k}:=\bm{\xi}_{0}^{k+1}$, $\mathbf{u}^{k}:=\mathbf{p}_{TD}^{k}$, and $\sigma^{k}:=e^{\lambda T^{k}}$, then the step-to-step discrete dynamics of the initial DCM by (4) is given by
$\mathbf{z}^{k}=\sigma^{k}\mathbf{z}^{k-1}-\mathbf{u}^{k},\quad k=1,2,...\quad,$ (8)
where $\mathbf{z}^{0}=\bm{\xi}_{0}^{1}=\bm{\xi}(0)$ is the initial DCM of the current step.
We now propose an MPC formulation based on the discrete dynamics (8) that minimizes the errors between the gait parameters $(\mathbf{z}^{k},\sigma^{k},\mathbf{u}^{k})$ and their desired values for multiple future steps. The design of the optimization constraints and objectives on the decision variables are as follows:
@@ III-1 Step Position
For the Digit robot model with underactuated ankles and flat feet, we define the step position $\mathbf{u}^{k}$ as the geometric center at the bottom of each stance foot. A stepping stone profile is given as a sequence of fixed positions in the world frame, while the robot states and variables such as the DCM and the step position are defined in each contact frame. Given the preview of next $N$ stepping stones in the current contact frame, the target is to minimize the error between the future step position $\mathbf{u}^{k}$ and stone position $\mathbf{p}_{stone}^{k}$ for $k=\{1,2,...,N\}$. Thus, the desired values of each step position are computed as
$\mathbf{u}_{\mathrm{des}}^{k}=\begin{cases}\mathbf{p}_{stone}^{1},\quad\quad\quad\quad\quad\ k=1\\ \mathbf{p}_{stone}^{k}-\mathbf{p}_{TD,\mathrm{ccf}}^{k-1},\quad k=2,...,N\end{cases},$ (9)
where $\mathbf{p}_{TD,\mathrm{ccf}}^{k-1}$ should be the actual step position of the future ($k$-1)-th step in the current contact frame. Since $\mathbf{p}_{TD,\mathrm{ccf}}^{k-1}$ are unknown during the current step, we use the previous planning results to approximate their actual values, which are iteratively updated inside the planner. In practice, if the planner has not returned the first result right after a touch-down, we then use the position of the ($k$-1)-th stone instead.
Moreover, the bound on each step position is defined as a rectangular area around the $p_{stone}^{k}$ that represents the physical dimension of each stepping stone.
@@ III-2 Step Duration
Since we emphasize the accuracy of the step position and viability of the walking gait, we make the step duration a slack variable to relax the problem. We choose a constant nominal value $T_{\mathrm{des}}^{k}=T_{\mathrm{nom}}$ and a constant bound $[T_{min},T_{max}]$ for $T^{k}$, and assign a small weight for the step duration term $\sigma^{k}$ in the cost function.
@@ III-3 Initial DCM
The desired value of the initial DCM $\mathbf{z}^{k}$ are computed using (5) by substituting the desired values of the step position and duration of the $k$-th step. The bounds of the initial DCM are obtained using the mechanical limits of the robot model, which rules out all the infeasible states that lead to the robot falling. However, these bounds alone do not guarantee foot placements that can yield viable states for all the given stepping stones. Therefore, we choose an appropriately high weight in the cost term for each $\mathbf{z}^{k}$ to track the desired value. This works as a soft constraint to enforce a viable evolution of the initial DCM through multiple steps.
Moreover, since this foot placement planner runs multiple times during a step, we use the measured value of the instantaneous DCM $\bm{\xi}_{\mathrm{mea}}(\tilde{t})$ to estimate the initial DCM of the current step $\mathbf{z}^{0}$ by computing it backward using (3):
$\mathbf{z}^{0}=\bm{\xi}_{0}^{1}=\bm{\xi}_{\mathrm{mea}}(\tilde{t})e^{-\lambda\tilde{t}},$ (10)
where $\tilde{t}\in[0,T^{1}]$ is the current relative time of the step so that $\mathbf{z}^{0}$ is also iteratively updated using the robot states.
@@ III-4 Discrete MPC Formulation
$\begin{split}\underset{\begin{subarray}{c}\mathbf{z}^{k},\sigma^{k},\mathbf{u}^{k}\end{subarray}}{\mathrm{arg\ min}}\quad\sum_{k=1}^{N}\beta^{k}\biggl{(}&\alpha_{z}^{k}\left\lVert\mathbf{z}^{k}-\mathbf{z}_{\mathrm{des}}^{k}\right\rVert^{2}\\ +&\alpha_{\sigma}^{k}|\sigma^{k}-\sigma_{\mathrm{des}}^{k}|^{2}\\ +&\alpha_{u}^{k}\left\lVert\mathbf{u}^{k}-\mathbf{u}_{\mathrm{des}}^{k}\right\rVert^{2}\biggr{)}\quad\\ \end{split}$
$\mathrm{s.\ t.}\quad$ $\mathbf{z}^{k}=\sigma^{k}\mathbf{z}^{k-1}-\mathbf{u}^{k},$
$\begin{bmatrix}z_{x,min}\\ z_{yl/r,min}\end{bmatrix}\leq\mathbf{z}^{k}\leq\begin{bmatrix}z_{x,max}\\ z_{yl/r,max}\end{bmatrix},$
$\quad\quad e^{\lambda T_{min}}\leq\sigma^{k}\leq e^{\lambda T_{max}},$
$\begin{bmatrix}u_{x,min}^{k}\\ u_{yl/r,min}^{k}\end{bmatrix}\leq\mathbf{u}^{k}\leq\begin{bmatrix}u_{x,max}^{k}\\ u_{yl/r,max}^{k}\end{bmatrix}.$
where $k=\{1,2,...,N\}$, and we use the weights $\alpha_{\square}^{k}$ and $\beta^{k}$ to represent soft constraints on the decision variable such that: $\alpha_{\square}^{k}$ assigns relative weights on $(\mathbf{z}^{k},\sigma^{k},\mathbf{u}^{k})$ for the $k$-th future step as we discussed above; $\beta^{k}$ is a decaying weight on each future step, addressing the importance of previewing earlier future steps, especially the imminent next one.
Since the discrete dynamics of DCM (8) takes a bilinear form about the decision variables $(\mathbf{z}^{k},\sigma^{k},\mathbf{u}^{k})$, we arrive at a nonlinear programming (NLP) problem. It takes considerably longer to solve than convex quadratic programming (QP), but it is still possible to implement in online foot placement planning thanks to this discrete formulation.
@@
This planner formulation is not limited to a specific type of terrain. Appropriate constraints and weights to represent different control objectives can be designed and implemented in a variety of flat-ground robotic walking scenarios.
@@ IV Adaptation of Variable Step Duration in the Low-level Controller
Figure 3: Illustration of updating $\tau(t)$ with respect to $T^{i}_{\mathrm{des}}$. For instance, when $T^{2}_{\mathrm{des}}$ is updated at $t^{2}$, $\tau(t)$ becomes the linear function that starts from $\tau^{2}$ and aims at $1$ if $t$ reaches $T^{2}_{\mathrm{des}}$.
The foot placement planner in Section III updates the desired step position $\mathbf{p}_{TD}^{1}$ and duration $T^{1}$ of the next step at a low frequency during the current swing phase, while the low-level controller aims to regulate the robot’s CoM dynamics towards the ALIP model and track a desired swing foot trajectory characterized by $(\mathbf{p}_{TD}^{1},T^{1})$ at a high frequency (e.g., $25$ Hz and $1$ kHz respectively). In this work, we use an existing whole-body controller presented in [17] to track these task space outputs and maintain a stable walking gait. The outputs of the low-level controller are defined using the robot configuration $q$ as
$\mathcal{Y}^{\mathrm{a}}(q):=\begin{pmatrix}\text{ base height }\\ \text{ torso orientation }\\ \text{ swing foot position }\\ \text{ swing foot orientation }\\ \end{pmatrix},$ (11)
where the desired position of the swing foot is generated as a 3-D polynomial trajectory ending at the desired step position with the desired step duration.
Note that the low-level controller uses a dimensionless phase variable $\tau$ to drive the trajectories of the robot, which increases monotonically from 00 to $1$ during the current swing phase. If the step duration is fixed, $\tau(t)$ is a linear function of the relative time $t$ with a fixed slope. However, when the desired step duration is variable during the swing phase, we need to compute the phase variable $\tau$ accordingly.
Let $n$ be the number of times the planner updates the desired foot placement during a swing phase, $t^{i}$ and $\tau^{i}$ ($i=\{0,1,...,n\}$) be the relative timestamp and the phase variable at the $i$-th planner update. As the swing phase starting from $t^{0}=0$ and $\tau^{0}=0$, the phase variable $\tau(t)$ is designed as:
$\tau(t)=\tau^{i}+(1-\tau^{i})\frac{t-t^{i}}{T_{\mathrm{des}}^{1,i}-t^{i}},\quad t\in[t^{i},t^{i+1}],$ (12)
where $T_{\mathrm{des}}^{1,i}$ is the $i$-th updated desired step duration. This design is primarily based on the following key points (See Fig. 3):
@@
    $\tau(t)$ increases monotonically from 00 to $1$ as $t$ increases from 00 to the last updated desired step duration $T_{\mathrm{des}}^{1,n}$;
@@
    $\tau(t)$ is continuous over $t$ and piecewise linear in all $[t^{i},t^{i+1}]$;
@@
    In each $[t^{i},t^{i+1}]$, $\tau(t)$ increases with a fixed slope such that it starts from $\tau^{i}$ and aims to reach $1$ if $t$ reaches $T_{\mathrm{des}}^{1,i}$, and then the slope is updated with the new $T_{\mathrm{des}}^{1,i+1}$ at $t=t^{i+1}$.
Moreover, the time derivative of the phase variable, $\dot{\tau}(t)$, is used to compute the time derivatives of the desired outputs $\dot{\mathcal{Y}}^{\mathrm{des}}$ inside the low-level controller, thus it can be computed using (12) as:
$\dot{\tau}(t)=\frac{1-\tau^{i}}{T_{\mathrm{des}}^{i}-t^{i}},\quad t\in[t^{i},t^{i+1}].$ (13)
@@ V Simulation Results
This foot placement planner is tested on the humanoid robot model Digit in the MuJoCo simulation environment [18]. The simulation runs at a fixed time step of $1$ ms, the same as the low-level controller. The foot placement planner runs every $40$ ms (i.e., 40 simulation time steps) and the nonlinear MPC in the foot placement planner is solved using the open-source solver IPOPT [19]. To balance the planner performance and computational efficiency, we choose the number of future steps considered in the MPC as $N=4$, which includes two complete cycles of 3-D bipedal walking. For all presented simulations, the parameters and constants of the foot placement planner are chosen as follows:
  • The step position bound is defined as a rectangular area of dimension $0.2\times 0.1$ meters centered at each $\mathbf{p}_{stone}^{k}$;
  • The step duration has a nominal value $T_{\mathrm{nom}}=0.5$ with the bound $[T_{min},T_{max}]=[0.35,0.65]$ seconds;
  • The initial DCM bounds are computed using the mechanical limits of the robot such as $\left\lvert x_{TD}^{k}\right\rvert\leq 0.6$ and $0.1\leq\left\lvert y_{TD}^{k}\right\rvert\leq 0.5$ meters with the step duration bound;
  • The cost weights for the $k$-th future step are chosen as $\alpha_{z}^{k},\alpha_{u}^{k}=[10,000,20,000]$, $\alpha_{\sigma}^{k}=1$, and $\beta^{k}=10^{4-k}$ for $k=\{1,2,3,4\}$, where we double the weights in the $y$-direction because the step position bound in $y$-direction is half of that in $x$-direction.
For a comprehensive test of the planner performance, four types of stepping stone profiles and a perturbation scenario with four different perturbations are designed to represent the challenging terrain on the flat ground, where we compare our $N$-step preview planner (denoted as ALIP-MPC) with the one-step preview planner in [10] (denoted as ALIP-QP) [^1] For animation video of the presented simulations, see https://youtu.be/DjH69m1kbnM..
@@
Profile No. $L^{j}$ (m) $W^{j}$ (m)
I fixed at $\{0.2,0.4,0.55\}$ fixed at $\{-0.15,0.15\}$
II $U[0.2,0.5]$ $U[-0.15,0.15]$
III (0.2(0.2( 0.2*$8,0.5$*8)8)8 )*$2$ $U[-0.15,0.15]$
IV $U[0.2,0.5]$ (0.1(-0.1( - $0.18,$ 0.1*8)8)8 )*$2$
[…]

### v3

@@ III Swing Foot Placement Planning via Discrete Model Predictive Control
The primary contribution of this work is to design a foot placement planner for stepping stones with high accuracy of step positions and enhanced resistance to external perturbations. The stepping stones can be modeled as a sequence of bounds on each step position, which leads to significant constraints on the dynamic walking. Therefore, multiple future steps must be considered to find a stable walking gait by exploiting the viable evolution of the initial DCM with variable step duration.
To simplify the denotation, we define $\mathbf{z}_{k}:=\bm{\xi}_{0,k+1}$, $\mathbf{u}_{k}:=\mathbf{p}_{TD,k}$, and $\sigma_{k}:=e^{\lambda T_{k}}$, then the step-to-step discrete dynamics of the initial DCM by (4) is given by
$\mathbf{z}_{k}=\sigma_{k}\mathbf{z}_{k-1}-\mathbf{u}_{k},\quad k=1,2,$$\quad$, (8)
where $\mathbf{z}_{0}=\bm{\xi}_{0,1}=\bm{\xi}(0)$ is the initial DCM of the current step for $k=1$.
We now propose a discrete MPC formulation based on the discrete dynamics (8) that minimizes the errors between the gait parameters $(\mathbf{z}_{k},\sigma_{k},\mathbf{u}_{k})$ and their desired values for multiple future steps. The design of the optimization constraints and objectives on the decision variables is as follows:
@@ III-1 Step Position
For the Digit robot model with underactuated ankles and flat feet, we define the step position $\mathbf{u}_{k}$ as the geometric center at the bottom of each stance foot. A stepping stone profile is given as a sequence of fixed positions in the world frame, while the robot states and variables, such as the DCM and the step position, are defined in the associated contact frame. Given the preview of next $N$ stepping stones in the current contact frame, the target is to minimize the error between each future step position $\mathbf{u}_{k}$ and its desired step position $\mathbf{u}_{\mathrm{des},k}$ for $k=\{1,2,...,N\}$, which is computed as
$\mathbf{u}_{\mathrm{des},k}=\begin{cases}\mathbf{p}_{stone,1},\quad\quad\quad\quad\quad\quad\ \ k=1\\ \mathbf{p}_{stone,k}-\mathbf{p}_{TDC,k-1},\quad k=2,...,N\end{cases}.$ (9)
Since each $\mathbf{u}_{k}$ is defined in its corresponding contact frame, $\mathbf{p}_{TDC,k-1}$ should be the actual step position of the future (kk-1)-th step in the current contact frame. Given that $\mathbf{p}_{TDC,k-1}$ are unknown during the current step, we use the previous planned values to approximate their actual values, which are iteratively updated inside the planner. In practice, if the planner has not returned the first result right after a touchdown, we then use the value of the (kk-1)-th stone position $\mathbf{p}_{stone,k-1}$ instead.
Moreover, the bound on each step position is defined as a rectangular region around each $\mathbf{p}_{stone,k}$ that represents the physical dimension of each stepping stone.
@@ III-2 Step Duration
Since we emphasize the accuracy of the step position and viability of the walking gait, we make the step duration a slack variable to relax the problem. We choose a constant nominal value $T_{k,\mathrm{des}}=T_{\mathrm{nom}}$ and a constant bound $T_{\mathrm{min}},T_{\mathrm{max}}]$ for $T_{k}$, and assign a small weight for the step duration term $\sigma_{k}$ in the MPC objective function.
@@ III-3 Initial DCM
The desired value of the initial DCM $\mathbf{z}_{k}$ is computed using (5) by substituting the desired values of the step position and duration of the $k$-th step. The bounds of the initial DCM are obtained using the mechanical limits of the robot model, which rules out all the infeasible states that lead to the robot falling. However, these bounds alone do not guarantee foot placements that can yield viable states for all the given stepping stones. Therefore, we choose an appropriately high weight in the MPC objective term for each $\mathbf{z}_{k}$ to track the desired value. This works as a soft constraint to enforce a viable evolution of the initial DCM through previewed stepping stones.
[…]
$\mathbf{z}_{0}=\bm{\xi}_{0,1}=\bm{\xi}_{\mathrm{mea}}(\tilde{t})e^{-\lambda\tilde{t}}$, (10)
where $\tilde{t}\in[0,T_{1}]$ is the current relative time of the step so that $\mathbf{z}_{0}$ is also iteratively updated using the robot states.
@@ III-4 Discrete MPC Formulation
[…]
$\mathrm{s.\ t.}\quad$ $\mathbf{z}_{k}=\sigma_{k}\mathbf{z}_{k-1}-\mathbf{u}_{k},$ (12)
$\begin{bmatrix}z^{x}_{\mathrm{min}}\\ z^{yl/r}_{\mathrm{min}}\end{bmatrix}\leq\mathbf{z}_{k}\leq\begin{bmatrix}z^{x}_{\mathrm{max}}\\ z^{yl/r}_{\mathrm{max}}\end{bmatrix}$, (13)
$\quad e^{\lambda T_{\mathrm{min}}}\leq\sigma_{k}\leq e^{\lambda T_{\mathrm{max}}},$ (14)
$\begin{bmatrix}u^{x}_{k,\mathrm{min}}\\ u^{yl/r}_{k,\mathrm{min}}\end{bmatrix}\leq\mathbf{u}_{k}\leq\begin{bmatrix}u^{x}_{k,\mathrm{max}}\\ u^{yl/r}_{k,\mathrm{max}}\end{bmatrix}$. (15)
where $k=\{1,2,...,N\}$, and we use the weights $\alpha_{(\cdot),k}$ (with subscripts $z,\sigma,u$) and $\beta_{k}$ to represent soft constraints on the decision variable such that: $\alpha_{(\cdot),k}$ assigns relative weights on $(\mathbf{z}_{k},\sigma_{k},\mathbf{u}_{k})$ for the $k$-th future step as we discussed above; $\beta_{k}$ is a decaying weight on each future step, addressing the importance of previewing earlier future steps, especially the imminent next one.
Although the discrete dynamics of DCM (8) takes a bilinear form about the decision variables $(\mathbf{z}_{k},\sigma_{k},\mathbf{u}_{k})$, such a discrete formulation significantly reduces the problem dimension compared to a continuous-time MPC. Furthermore, we can utilize the latest commercial solvers to achieve online implementation.
@@
This MPC formulation is not limited to a particular terrain type. Appropriate weights and bounds to represent diverse control objectives, such as more accurate step position or more resistance to perturbation, can be designed and implemented in a variety of flat-ground robotic walking scenarios. Notably, both objectives can be interpreted as equivalent constraints on the evolution of the DCM.
@@ IV Adaptation of Variable Step Duration in the Low-level Controller
Figure 3: Illustration of updating $\tau(t)$ with respect to $(T_{1,\mathrm{tgt}})_{i}$. For instance, when $(T_{1,\mathrm{tgt}})_{2}$ is updated at $t_{2}$, $\tau(t)$ transitions linearly from $\tau_{2}$ to 11 if $t$ reaches $(T_{1,\mathrm{tgt}})_{2}$, until $t$ reaches $t_{3}$ when the 3rd update takes place.
The foot placement planner in Section III updates the target step position $\mathbf{p}_{TD,1,\mathrm{tgt}}$ and duration $T_{1,\mathrm{tgt}}$ of the current step at a low frequency during the current swing phase, while the low-level controller aims to regulate the robot’s CoM dynamics towards the ALIP model and track a desired swing foot trajectory characterized by $(\mathbf{p}_{TD,1,\mathrm{tgt}},T_{1,\mathrm{tgt}})$ at a high frequency. In this work, we use an existing whole-body controller presented in [1] to track these task space outputs and maintain a stable walking gait. The outputs of the low-level controller are defined using the robot configuration $q$ as
[…]
where the desired swing foot trajectory is generated as a 3-D polynomial curve ending at the desired step position with the desired step duration.
Note that the low-level controller uses a dimensionless phase variable $\tau$ to drive the trajectories of the robot, which increases monotonically from 0 to $1$ during the current swing phase. If the desired step duration is constant, $\tau(t)$ is a linear function of the relative time $t$ with a fixed slope. However, when the desired step duration is variable during the swing phase, the phase variable $\tau$ needs to be computed accordingly to reflect the time-varying trajectory.
Let $n$ be the number of times the planner updates the desired foot placement during a swing phase, $t_{i}$ and $\tau_{i}$ ($i=\{0,1,...,n\}$) be the relative timestamp and the phase variable at the $i$-th planner update. As the swing phase starting from $t_{0}=0$ and $\tau_{0}=0$, the phase variable $\tau(t)$ is computed as:
$\tau(t)=\tau_{i}+(1-\tau_{i})\frac{t-t_{i}}{(T_{1,\mathrm{tgt}})_{i}-t_{i}},\quad t\in[t_{i},t_{i+1}],$ (17)
where $(T_{1,\mathrm{tgt}})_{i}$ is the $i$-th updated tgtired step duration. This design is primarily based on the following key points (See Fig. 3):
@@
    $\tau(t)$ increases monotonically from 0 to $1$ as $t$ increases from 0 to the last updated desired step duration $(T_{1,\mathrm{tgt}})_{n}$;
@@
    $\tau(t)$ is continuous over $t$ and piecewise linear in all $t_{i},t_{i+1}]$;
@@
    In each $t_{i},t_{i+1}]$, $\tau(t)$ increases with a fixed slope such that it starts from $\tau_{i}$ and aims to reach $1$ if $t$ reaches $(T_{1,\mathrm{tgt}})_{i}$, until the slope is updated with the new $(T_{1,\mathrm{tgt}})_{i+1}$ at $t=t_{i+1}$.
Moreover, the time derivative of the phase variable $\dot{\tau}(t)$ is used to compute the time derivatives of the desired outputs $\dot{\mathcal{Y}}_{\mathrm{tgt}}(q)$ inside the low-level controller, which can be computed from (17) as:
$\dot{\tau}(t)=\frac{1-\tau_{i}}{(T_{1,\mathrm{tgt}})_{i}-t^{i}},\quad t\in[t_{i},t_{i+1}].$ (18)
@@ V Simulation Results
The proposed foot placement planner is tested on the humanoid robot model Digit in the MuJoCo simulation environment [15] on a personal laptop with an Intel i7 processor. The simulation runs at a fixed timestep of $1$ ms, the same as the low-level controller. The foot placement planner runs every $20$ ms, where the nonlinear MPC is solved by a state-of-the-art solver FORCES PRO [18] (with average solving time around 5 ms). The number of future steps previewed in the MPC is chosen as $N=4$ to include two complete cycles of 3-D bipedal walking, though values up to 8 do not significantly increase the computation time. For all presented simulations, the parameters and constants of the foot placement planner are chosen as follows:
  • The MPC objective emphasizes enforcing the nominal initial DCM with accurate step position, thus the weights for the $k$-th future step are chosen as $(\alpha_{z,k},\alpha_{\sigma,k},\alpha_{u,k})=(1\mathrm{e}4,1,2\mathrm{e}4)$, and $\beta_{k}=10^{4-k}$ for $k=\{1,2,3,4\}$, where we double the weights in the $y$-direction because the step position bound in $y$-direction is half of that in $x$-direction;
  • The step position bound is defined as a $0.2\times 0.1$ meters rectangular region centered at each $\mathbf{p}_{stone,k}$;
  • The step duration has a nominal value $T_{\mathrm{nom}}=0.5$ with the bound $T_{\mathrm{min}},T_{\mathrm{max}}]=[0.3,0.7]$ seconds;
  • The initial DCM bounds are computed using the mechanical limits of the robot such as $|p^{x}_{TD,k}|\leq 0.6$ and $0.1\leq|p^{y}_{TD,k}|\leq 0.5$ meters with the step duration bound above.
For a comprehensive test of the planner performance, four types of stepping stone profiles and a perturbation scenario with four different perturbations are designed to represent the challenging terrain on flat ground, where we compare our $N$-step preview planner (denoted as DCM-MPC) with the one-step preview planner in [8] (denoted as DCM-QP).
@@
Profile No. $L_{j}$ (m) $W_{j}$ (m)
I fixed at $\{0.2,0.4,0.6\}$ fixed at $\{-0.15,0.15\}$
II $U[0.2,0.5]$ U[0.15,0.15]U[-0.15,0.15]
III (0.2(0.2*8,0.58,0.5*8)8)*22 U[0.15,0.15]U[-0.15,0.15]
IV $U[0.2,0.5]$ (0.1(-0.1*8,0.18,0.1*8)8)*22
  • Profile I contains $3\times 2$ combinations of fixed $L_{j}$ and $W_{j}$; Profile $I\simIV$ includes one or two lists of uniformly distributed random values as $L_{j}$ and/or $W_{j}$, denoted as $U[\cdot,\cdot]$; Profile III uses a list of alternating values as $L_{j}$, which is 8 consecutive values of $0.2$ followed by 8 values of $0.5$ and repeated once, which emphasizes on abrupt changes of stone positions in the $x$-direction (similar to $W_{j}$ in Profile IV).